OpenStudy (anonymous):

Check my answer plz If a stone is tossed from the top of a 230 meter building, the height of the stone as a function of time is given by h(t) = -9.8t^2 – 10t + 230, where t is in seconds, and height is in meters. After how many seconds will the stone hit the ground? Round to the nearest hundredth’s place; include units in your answer

OpenStudy (anonymous):

h(t) = -9.8t2 – 10t + 230 -230 = -9.8t2 – 10t + 230 9.8 t^2 + 10 t - 460 =0 4.9 t^2 + 5 t - 230 =0 t = { -5 + sq rt [ 25 +4508] } / 9.8 = 6.36 s ---answer t1 = { -5 - sq rt [ 25 +4508] } / 9.8 not valid

OpenStudy (anonymous):

you made a mistake here but not in your calculation

OpenStudy (anonymous):

the fact that the building is 230 meters tall is already taken care of in the equation. that is why it has the 230 as the "constant" at teh end

OpenStudy (anonymous):

your calculation is correct, but you should just take the original equation and set it = 0 and solve in other words solve \[ -9.8t^2 – 10t + 230=0\]

OpenStudy (anonymous):

k

OpenStudy (anonymous):

h(t) = -9.8t^2 – 10t + 230 =0 ==> use dicriminant rule then

OpenStudy (anonymous):

in other words you were thinking too hard. the equation actually gives the height and you don't have to adjust

OpenStudy (anonymous):

beacuse when the stone hits the ground ==> h=0

OpenStudy (anonymous):

so if you use quadratic formula you should get two values for t, one negative an done positive. The positive value is your answer and should be approximately 4.36 sec.

OpenStudy (anonymous):

7(x-4)<4(2x-1)